OpenStudy (anonymous):

what is the factored form of d2 + 12d + 32 ?

OpenStudy (anonymous):

(d+8)(d+4)

OpenStudy (anonymous):

\[d^2 +12d+32\]is this how the equation looks or what you mean?

OpenStudy (ajprincess):

First we need to find two numbers whose sum is 8 and product is 32. 8 and 4 are the two numbers that satisfy this. Then d^2+8d+4d+32 =d(d+8)+4(d+8) =(d+8)(d+4)